Hi, I am having trouble with this exercise:
11sin^2y = 13-3sin^2y for y, where 0<= y <360, How many angles satisfy this equation?

Answer by stanbon(48558)

11sin^2y = 13-3sin^2y for y, where 0<= y <360, How many angles satisfy this equation?
--------------
14sin^2y=13
sin^2y=13/14=0.92857...
siny=0.963624
y=arcsin(0.963624..)=74.498.. degrees or 105.501.. degrees
Cheers,
Stan H.